OpenStudy (anonymous):

Which of the following is a solution to the equation x 2 – 6x + 5 = 0?

OpenStudy (radar):

\[x ^{2}-6x+5=0\] factor getting (x-5)(x-1)=0. If a product of factors is 0 (as above) then the factors are 0 or at least one of them!. Set the factors to 0 and solve. x-5=0, x=5 x-1=0, x=1